OpenStudy (surryyy):

The area of a triangle is 150 square feet. If the base is x feet and the height is x + 5 feet, calculate the height. Answer 15 20 10 5

OpenStudy (surryyy):

my answer is 10, is it correct?

OpenStudy (anonymous):

A=1/2b*h 150=1/2*x*(x+5_) 300=x*(x+5) 300=x^2+5x 300=(x-15)(x+20) x=15 or x=-20 since area cannot be negative your answer is x=15

OpenStudy (surryyy):

oh ok thnx!

OpenStudy (surryyy):

The correct answer was 20.